# From Text to Talk: Audio-Language Model Needs Non-Autoregressive Joint Training

### v2

@@ A.3 Block-wise Masked Diffusion Generation For Audio Tokens
A.4 Training dataset details
@@
A.5 Evaluation Deatils
A.5.1 Evaluation Tasks
@@
Previous evaluation frameworks such as Kimi-Audio assess Audio-QA performance using the text portion of interleaved outputs, which overlooks the fact that the audio output of an end-to-end speech model more directly reflects its ability to generate natural and semantically faithful responses. To address this limitation, we evaluate Audio-QA directly on the audio outputs of our framework by first applying an ASR model to transcribe the generated audio into text, where Whisper-Large-v3 is used for English audio and Paraformer-zh for Chinese audio, with a comparison of ASR performance across different models provided in Table 3. The transcribed text is then combined with the original QA queries and the ground truth answers and passed to a large scale reasoning model, Qwen3-235B-A30B, which serves as an LLM-as-a-Judge model to determine whether the response semantically matches the reference and to provide either a correctness label or a graded score. We report the average accuracy or score on the benchmark, and this evaluation pipeline provides a more faithful assessment of our model’s audio-to-audio QA ability in realistic conversational scenarios where speech serves as the output modality.
@@
ASR Task
A.5.2 Evaluation Datasets
@@
We evaluate model performance on a diverse set of benchmarks covering both Audio Question Answering (Audio-QA) and Automatic Speech Recognition (ASR) tasks. For Audio-QA, we use four datasets: AlpacaEval, TriviaQA, and WebQuestions (English), along with ReasoningQA (Chinese), assessing cross-lingual reasoning and comprehension from speech. For ASR, we include five datasets: Fleurs-zh/en (multilingual), AISHELL-1/2, and WenetSpeech (all Chinese), covering varied domains, accents, and recording conditions to robustly measure transcription accuracy. Dataset details are summarized in Table 4.
Dataset Language Task Type
AlpacaEval English Audio-QA
ReasoningQA Chinese Audio-QA
TriviaQA English Audio-QA
WebQuestions English Audio-QA
Fleurs-zh Chinese ASR
AISHELL-2 Chinese ASR
AISHELL-1 Chinese ASR
WenetSpeech Chinese ASR
Fleurs-en English ASR
Table 4: Evaluation datasets used for Audio-QA and ASR tasks.
A.6 Data Format of Training Data

### v3

@@ A.4 Block-wise Masked Diffusion Generation For Audio Tokens
A.5 Illustration of Training Loss and Attention Design
Figure 3(a) shows the training loss of our framework: AR loss is applied to text spans, while NAR loss—based on discrete diffusion—is used for audio spans. Although we employ discrete diffusion, our framework is extensible to other NAR generation methods. Figure 3(b) visualizes the attention pattern described in Section 3.4.
A.6 Dataset details
Table 4 provides a summary of the training datasets, with detailed examples provided in the Appendix A.9.1. During training, we aim to construct a balanced corpus that supports effective learning across multiple tasks. Specifically, we randomly sample one million instances from the ASR dataset, the TTS dataset, and the audio chat dataset respectively. In addition, we create bilingual interleaved text and audio data, ensuring that Chinese and English are represented in approximately equal proportions. To build the audio chat corpus, we rely on the text-to-audio dataset VoiceAssistant-400K together with the text-based datasets OpenHermes-2.5 and Firefly-Train-1.1M, and we employ a TTS model, namely CosyVoice2, to convert text into synthetic audio so as to enrich the training data. To further enhance cross-modal alignment between text and audio, we follow prior work (Zeng et al., 2025) and supplement the training corpus with interleaved text and audio data derived from the large-scale pretrained corpus FineWeb-Edu. This strategy not only expands task coverage but also strengthens the model’s ability to jointly learn from and align textual and acoustic modalities. The evaluation datasets are shown in Table 7.
@@ A.6 Dataset details
A.7 Evaluation Deatils
A.7.1 Evaluation Tasks
URO-Bench
We leverage URO-Bench Yan et al. (2025) for a more comprehensive evaluation of our proposed method against existing baselines. URO-Bench is specifically designed for audio-in, audio-out tasks, directly simulating real-world conversational scenarios. In this framework, the spoken outputs from each model are first transcribed into text using Whisper-large Radford et al. (2022), and the resulting transcripts are then evaluated for correctness, coherence, and task alignment. This evaluation employs a hybrid scoring framework comprising three components: (1) an LLM-as-a-judge (originally implemented via commercial LLM APIs) to assess semantic correctness and task alignment; (2) rule-based metrics for automatic Word Error Rate (WER) computation; and (3) a fine-tuned emotion-aware model to evaluate the appropriateness of affective expression in spoken responses. Together, these components ensure that model outputs are judged not merely on surface-level textual fidelity, but on semantic accuracy, transcription quality, and emotional coherence.
URO-Bench structures its evaluation across two difficulty levels: Basic and Pro. Each level comprises three distinct categories: Understanding tasks, Reasoning tasks, and Oral Conversation tasks. This hierarchical design facilitates a fine-grained assessment of model capabilities across increasing levels of linguistic and cognitive demand, covering both single-round and multi-round scenarios. In our experiments, we report results on the English subset of URO-Bench’s evaluation set. Due to limited access to the original evaluation APIs (Gemini Flash and GPT-4o-mini), we substitute the LLM-as-a-judge component with thinking model Qwen3-30B-A3B, while keeping all other components—including the Whisper-based ASR pipeline and rule-based scoring—identical to the original implementation.
Furthermore, the benchmark integrates a perceptual quality evaluation mechanism. We employ the strong UTMOS Saeki et al. (2022) for the UTMOS score and DNSMOS Dubey et al. (2024) for the NMOS score evaluation, enabling the joint assessment of content accuracy and acoustic quality. Importantly, none of the URO-Bench data was used during training or validation, ensuring an unbiased assessment of generalization.
@@ A.7.1 Evaluation Tasks
In addition to URO-Bench, we also evaluate our model with the Audio-QA tasks established by Kimi-Audio Ding et al. (2025). Previous evaluation framework in Kimi-Audio assess Audio-QA performance using the text portion of interleaved outputs, which overlooks the fact that the audio output of an end-to-end speech model more directly reflects its ability to generate natural and semantically faithful responses. To address this limitation, we evaluate Audio-QA directly on the audio outputs of our framework by first applying an ASR model to transcribe the generated audio into text, where Whisper-Large-v3 Radford et al. (2022) is used for English audio and Paraformer-zh for Chinese audio, with a comparison of ASR performance across different models provided in Table 5. The transcribed text is then combined with the original QA queries and the ground truth answers and passed to a large scale reasoning model, Qwen3-235B-A30B, which serves as an LLM-as-a-Judge model to determine whether the response semantically matches the reference and to provide either a correctness label or a graded score. We report the average accuracy or score on the benchmark, and this evaluation pipeline provides a more faithful assessment of our model’s audio-to-audio QA ability in realistic conversational scenarios where speech serves as the output modality.
ASR Task
AAC Task
To assess the model’s capacity to comprehend complex or acoustically challenging audios, we evaluate its audio captioning (AAC) performance on two established benchmarks: Clotho-v2 Drossos et al. (2020) and MACS Martín-Morató and Mesaros (2021). We input audio clips and generate corresponding textual captions. The quality of these captions is then evaluated using Qwen3-30B-A3B Yang et al. (2025a), guided by the evaluation prompt introduced in CLAIR-A Wu et al. (2024), which emphasizes semantic relevance, completeness, and naturalness. The judge assigns a score on a 0–100 scale, where higher scores indicate better caption quality.
@@
A.7.2 Evaluation Datasets
URO-Bench
We use the English portion of URO-Bench Yan et al. (2025) to evaluate our model’s performance. As detailed in Table 6, the benchmark consists of 10 basic tasks and 12 pro tasks. The basic tasks include 4 oral conversation, 4 reasoning, and 2 understanding tasks, while the pro tasks comprise 4 understanding, 4 reasoning, and 4 oral conversation tasks. The final score is obtained by first averaging the model’s performance on each dataset, and then averaging these scores within each (difficulty, category) group.
Table 6: Evaluation datasets used from URO-Bench.
Dataset Task /Evaluation Aspect data nums Category
Basic tasks
AlpacaEval Authentic, open-ended dialogue 199 Oral Conversation
CommonEval Authentic, open-ended dialogue 200 Oral Conversation
WildchatEval Real-world conversation 349 Oral Conversation
StoralEval Deduce morals from a given story 201 Reasoning
Summary Summarize a given story or statement 118 Oral Understanding
TruthfulEval Factual questions about life 470 Reasoning
GaokaoEval English listening questions 303 Understanding
Gsm8kEval Practical mathematical problems 582 Reasoning
MLC Mathematics, logic, and common sense 177 Reasoning
Repeat Repeat the user’s words verbatim 252 Understanding
Pro Tasks
CodeSwitching-en Understand code switching sentences 70 Understanding
GenEmotion-en Respond in a specified tone 54 Oral Conversation
GenStyle-en Respond in a specified style 44 Oral Conversation
MLCpro Difficult mathematical, scientific questions 91 Reasoning
Safety-en Reject answering privacy-related questions 24 Reasoning
SRT-en Sing, recite poems, read tongue twisters 43 Oral Conversation
UnderEmotion-en Understand the speaker’s mood 137 Understanding
Multilingual Respond in multiple languages 1108 Oral Conversation
ClothoEval-en Comprehension of general ambient sounds 265 Understanding
MuChoEval-en Comprehension of music 311 Understanding
MtBenchEval-en Multi-round spoken dialogue 190 Reasoning
SpeakerAware-en Multi-speaker multi-round dialogues 55 Reasoning
Audio-QA, ASR and AAC Task
@@ A.7.2 Evaluation Datasets
We evaluate model performance on a diverse set of benchmarks covering both Audio Question Answering (Audio-QA), Automatic Speech Recognition (ASR), and automatic audio caption (AAC) tasks. For Audio-QA, we use four datasets: AlpacaEval, TriviaQA, and WebQuestions (English), along with LLaMAQuestions (English), assessing cross-lingual reasoning and comprehension from speech. For ASR, we include five datasets: Fleurs-zh/en (multilingual), AISHELL-1/2, and WenetSpeech (all Chinese), covering varied domains, accents, and recording conditions to robustly measure transcription accuracy. For AAC, we include two datasets: Clotho-v2 (English) and MACS (English), covering natural audios collected from various environmental sound clips, not limited to human-to-human dialogue, which helps assess the model’s ability to understand the environment rather than simple language processing. Dataset details are summarized in Table 7.
Table 7: Evaluation datasets used for Audio-QA, ASR and AAC tasks.
Dataset Language Task Type Abbreviation
AlpacaEval English Audio-QA AE.
LLaMAQuestions English Audio-QA LQ.
TriviaQA English Audio-QA TQA.
WebQuestions English Audio-QA WQ.
Fleurs-zh Chinese ASR Fzh.
AISHELL-2 Chinese ASR A2.
AISHELL-1 Chinese ASR A1.
WenetSpeech-test_meeting Chinese ASR WS_m.
WenetSpeech-test_net Chinese ASR WS_n.
Fleurs-en English ASR Fen.
Clotho-v2 English AAC Clo.
MACS English AAC MACS
A.8 Baselines
We compare our TtT model with the following state-of-the-art large audio-language models to evaluate its effectiveness:
  • Moshi Défossez et al. (2024): It unifies streaming speech and text understanding within a single autoregressive framework, aligning acoustic and linguistic representations for low-latency real-time dialogue and robust multimodal instruction following.
  • SpeechGPT Zhang et al. (2023): It incorporates discrete speech tokens into a single language model and follows a three-stage training pipeline to enable unified speech–text understanding and cross-modal instruction following within one framework.
  • Kimi-Audio Ding et al. (2025): It uses a multi-task training pipeline to align speech, text and semantics through contrastive and generative objectives, enabling robust instruction-following and long-form audio dialogue understanding.
  • VITA-Audio Long et al. (2025): It tackles the latency bottleneck in LSLMs by introducing a fast interleaved decoding mechanism and dynamic token predictor, allowing efficient and streaming-capable audio response generation.
  • LLaMA-Omni Fang et al. (2025): It extends a unified language model to support real-time speech understanding and generation by integrating low-latency audio streaming, codec-based tokenization, and tightly aligned speech–text representations for seamless multimodal interaction.
  • GLM-4-Voice Zeng et al. (2024): It introduces a unified end-to-end spoken language model that interleaves speech and text modalities using a supervised speech tokenizer and joint training paradigm, enabling high-quality spoken dialogue generation.
  • Mini-Omni Xie and Wu (2024a): It enables real-time speech interaction by generating text and audio tokens in parallel within one model, using text-instructed parallel decoding and a lightweight training pipeline to preserve the base model’s reasoning ability.
  • SLAM-Omni Chen et al. (2025): It enables end-to-end spoken dialogue by modeling text and semantic audio tokens in parallel within a single model, supporting zero-shot timbre control and low-latency voice interaction through single-stage training.
A.9 Training Details
A.9.1 Data Format of Training Data
@@
A.10 Usage of LLM
In this paper, the LLM is employed solely for text refinement—correcting typos, fixing spelling errors, and enhancing readability. It is not used for generating research ideas, producing results, or creating original content.